OpenStudy (anonymous):

Solve the equation in the interval [0,2pi]. 2sin(t)cos(t)-cos(t)+2sin(t)-1=0. I already factored it and my answers were pi/2, 2pi/3, 4pi/3 but I got it wrong.

OpenStudy (freckles):

you factored by grouping , correct?

OpenStudy (anonymous):

Yass

OpenStudy (freckles):

cos(t)[2 sin(t)-1]+1[2 sin(t)-1]=0 [2 sin(t)-1](cos(t)+1]=0

OpenStudy (freckles):

then you set two two factors equal to zero?

OpenStudy (freckles):

are these the equations you got 2sin(t)-1=0 cos(t)+1=0 ?

OpenStudy (anonymous):

No I got cos(t)=0 and sin(t)-1=0

OpenStudy (freckles):

do you see how I got my equations?

OpenStudy (anonymous):

yeah i did thank you so much cx

OpenStudy (freckles):

 drew it just in case it made it more prettier to you